OpenStudy (anonymous):

Please help! will give medal :) Write the expression as the sine, cosine, or tangent of an angle. Cos 107 degrees cos 42 degrees + sin 107 degrees sin 42 degrees

OpenStudy (solomonzelman):

I am not sure what you really need to do here.

OpenStudy (anonymous):

Idk thats what the instructions are, choices are a. sin 65, b. cos 65, c. cos 149, d. sin 149

OpenStudy (solomonzelman):

I am thinking of sin(a+b) identities right now....

OpenStudy (solomonzelman):

Oh, lets work this backwards, look at this formula below, cos(A-B)=cos A cos B + sin A sin B and compare to, cos(65)=cos(107-42)=cos (107) cos (42) + sin (107) sin (42)

OpenStudy (solomonzelman):

reply when you figure, it shouldn't be hard ;)

OpenStudy (anonymous):

thanks, I think I got that one, but theres something I'm missing with the concept because I can't do the next problem either which is similar... same instructions but sin 5 x cos x - cos 5x sin x

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

DO you have the options?

OpenStudy (anonymous):

a. cos 6x b. cos 4x c. sin 6x d. sin 4x

OpenStudy (solomonzelman):

alright, look at this formula. sin(A-B)=sin A cos B - cos A sin B and compare to, sin(5x)=sin(6x-x )=sin (5x) cos (x) - cos (5x) sin (x)

OpenStudy (anonymous):

so sin 6x?

OpenStudy (solomonzelman):

I am getting sin 5x

OpenStudy (anonymous):

thats not an option :O

OpenStudy (solomonzelman):

Yeah, ik, I am thinking...

OpenStudy (solomonzelman):

are you sure you typed the question correctly?

OpenStudy (anonymous):

sin 5x cos x - cos 5x sin x

OpenStudy (solomonzelman):

then it's sin(4x)=sin(5x-x )=sin (5x) cos (x) - cos (5x) sin (x) my bad abt the 1st mistake :(

OpenStudy (anonymous):

ohh, all good, thanks!

OpenStudy (solomonzelman):

You welcome, have a good one ;)